OpenStudy (anonymous):

Im struggling with modulo arithmetic. How would i calculate 3^100 mod 1000

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \normalsize \text{the shortest way would be to use binomial theorm} \hspace{.33em}\\~\\ 3^{100} \pmod {1000} \hspace{.33em}\\~\\ =9^{50} \pmod {1000} \hspace{.33em}\\~\\ =(-1+10)^{50} \pmod {1000} \hspace{.33em}\\~\\ =\dbinom{50}{0}\cdot (-1)^{50}\cdot (10)^{0}+\dbinom{50}{1}\cdot (-1)^{49}\cdot (10)^{1}\hspace{.33em}\\~\\ +\dbinom{50}{2}\cdot (-1)^{48}\cdot (10)^{2}+ \color{red}{\dbinom{50}{3}\cdot (-1)^{47}\cdot (10)^{3}\cdot \cdot \cdot } \pmod {1000} \hspace{.33em}\\~\\ \normalsize \text{note :-(starting from red terms all will be) } \equiv 0 \pmod {1000} \hspace{.33em}\\~\\ =1+50\cdot (-1)\cdot (10)^{1} +\dfrac{50\times 49}{2}\cdot 1\cdot 100 \hspace{.33em}\\~\\ + \color{red}{0+0+0\cdot \cdot \cdot } \pmod {1000} \hspace{.33em}\\~\\ =1-500 +2500\times 49 \pmod {1000} \hspace{.33em}\\~\\ =1-500 +500\times (50-1) \pmod {1000} \hspace{.33em}\\~\\ =1-500 +25000-500 \pmod {1000} \hspace{.33em}\\~\\ =1 \pmod {1000} \hspace{.33em}\\~\\ =1 \hspace{.33em}\\~\\ }\end{align}\)